Suppose the following data describe output in two different years: Item Apples Bicycles Movie rentals Year 1 20,000 @ 25c eaclh 700 @ $800 each Year 2 30,000 @ 30c eaclh 650 @ $900 each 12,000 @ $1.50 each 10,000 $1.00 each a. Compute nominal GDP in each year. Year 1: Year 2: b. By what percentage did nominal GDP change between Year 1 and Year 2? (Use the percentage change formula [(new value - original value)/original valuel*100.) Nominal GDP by c. Now compute real GDP in Year 2 by using the prices of Year 1 Real GDP: d. By what percentage did real GDP increase between Year 1 and Year 2? (Use the percentage change formula - [(new value - original value)/original valuel* 100.) Real GDP by
Answer #1

Answer a : Nominal GDP in year 1 : (20,000* $0.25+ 700*$800+ 10,000*$1)= $5000+ $560000+ $10,000= $5,75,000

Nominal GDP in year 2 = ( 30,000*$0.30+ 650*$900+ 12000*$1.50)= ( $9000+ $585000+$18000)= $612000

b : Nominal GDP % change = ( (New value - Original value) / original value)*100 = ((6,12,000- 5,75,000)/5,75,000)*100= 6.43%

Nominal GDP has been change from $575000 to $612000.

Nominal GDP by 6.43%

c : Real GDP in year 2 is = price of base year * Quantity of current year = (30,000* $0.25+ 650*$800+ 12000*$1)= $7500+ $5,20,000+ $12000)= $5,39,500

d : Real GDP change = ( ($539,500- $575000)/5,75,000)*100

Real GDP % change = -6.17%

Real GDP has been decreased.

Real GDP has been $5,75,000 to $5,39,500